OpenStudy (anonymous):

If the given angle is in standard position, find two positive coterminal angles and two negative coterminal angles. (Enter your answers as a comma-separated list.) (a) 700° positive angles negative angles (b) 5π/7 positive angles negative angles (c) −π/3 positive angles negative angles

OpenStudy (solomonzelman):

Do you know what a coterminal angle is?

OpenStudy (solomonzelman):

I was trying to start a discussion and help you, but you left....

OpenStudy (anonymous):

Sorry! i'm here

OpenStudy (anonymous):

Are you still here?

OpenStudy (anonymous):

Sorry i was finding out how to use this site. This is my first time on it.

OpenStudy (solomonzelman):

I have to go now tho.... (I am going to be back in an hour). I'll post some instructions for you . @Reaper534 maybe can help you. Sorry for bugging you, reaper..

OpenStudy (anonymous):

for sure

OpenStudy (anonymous):

now lets see

OpenStudy (anonymous):

I dont think ive done this yet not sure how its solved

OpenStudy (anonymous):

@AriPotta Do you know this?

OpenStudy (solomonzelman):

I am back. (I was studying my Jewish stuff with a group...) Now, Coterminal angle: For angle (lets say) 60, it will be -300, -660, ... (subtracting 360 each time) 420, 780 ... (adding 360 each time)

OpenStudy (anonymous):

hey i got it figured out! thank you so much though

OpenStudy (solomonzelman):

The idea is, 30 and 390 is the same exact angle, see?

OpenStudy (anonymous):

I apologize if i wasted your time.

OpenStudy (anonymous):

That actually helps me understand it a bit better though. Thank you (:

OpenStudy (solomonzelman):

I am sorry, I really had to go.... I should have given you the instructions or advices, but my connection snapped. Just like many times when I try to put some fancy codes.

OpenStudy (anonymous):

Its okay! you actually explained it very well, someone else helped me out dont worry. Can i fan you and if i have another question ill ask?

OpenStudy (solomonzelman):

You can ask me without fanning me. Whenever I am online. The thing is tho, that I am rarely online, due to my college.

OpenStudy (anonymous):

Okay, if you're not busy right now i have another question..

OpenStudy (solomonzelman):

Sure, go ahead.

OpenStudy (anonymous):

okay, thank you so much, one second

OpenStudy (anonymous):

Approximate, to the nearest 0.01 radian, all angles θ in the interval [0, 2π) that satisfy the equation. (Enter your answers as a comma-separated list.) (a) sin θ = −0.0134 θ = (b) cos θ = 0.9238 θ = (c) tan θ = 0.46 θ = (d) cot θ = −2.771 θ = (e) sec θ = −3.5 θ = (f) csc θ = 1.22 θ =

OpenStudy (anonymous):

If it shows the question marks, try backing out and coming back into the conversation

OpenStudy (solomonzelman):

you can refresh, and this is what I'll do.

OpenStudy (solomonzelman):

I don't have a calculator with me, but lets say for the first problem. Can you take the inverse sine? sin^(-1) −0.0134 = find this, and then convert to radians.

OpenStudy (anonymous):

Oh okay that makes sense.

OpenStudy (anonymous):

but how do i take the inverse of the ones that arnt sin cos or tan

OpenStudy (anonymous):

same thing?

OpenStudy (solomonzelman):

well, I am sure there is something like inverse cot, sec and csc. You CAN though, re-write them, using the rules below: csc^-1 x = 1 / sin^-1 x sec^-1 x = 1 / cos^-1 x cot^-1 x = 1 / tan-1 x It is the same thing for the inverses of sin cos and tan .

OpenStudy (solomonzelman):

you can take an inverse secant (for instance) . (Nothing wrong with that)